I am having trouble setting up this problem: In Triangle ABC, the measure of angle A is 141 degree's more than the measure of angle B. The measure of angle B is the same as the measure of angle C. What is the measure of each angle.

Found 2 solutions by checkley77, ankor@dixie-net.com:
Answer by checkley77(12569)

B=C
A=B+141 OR A-B=141
A+B+C=180
A+2B=180
-(A-B=141) SUBTRACT THESE TWO EQUATIONS
----------------------
2B+B=39
3B=39
B=39/3
B=13 DEGREES. C=13 DEGREES.
A=13+141
A=154 DEGREES.
PROOF:
154+13+13=180
180=180

Answer by ankor@dixie-net.com(15649)

I am having trouble setting up this problem: In Triangle ABC, the measure of
angle A is 141 degree's more than the measure of angle B.
The measure of angle B is the same as the measure of angle C.
What is the measure of each angle.
:
Let x = angle B
then
(x+141) = angle A; " A is 141 degree's more than the measure of angle B."
x = C; "measure of angle B is the same as the measure of angle C."
:
A + B + C = 180
(x+141) + x + x = 180
:
Find x and you have angle B